# Propagation of Chaos for Mean-Field Langevin Dynamics and its Application to Model Ensemble

### v1

@@ 2.2 Mean-field Langevin dynamics and uniform-in-time propagation of chaos
and $L_{2}$-regularizaton $r(x)=\lambda^{\prime}\|x\|_{2}^{2}$. We assume that $\sup_{x\in\mathbb{R}^{d},z\in\mathcal{Z}}|h(x,z)|\leq R$ and that for any $y\in\mathbb{R}$, $\ell(\cdot,y)$ is convex and $L$-smooth; there exists $L>0$ such that for any $a,b\in\mathbb{R}$, $\ell(b,y)\leq\ell(a,y)+\frac{\partial\ell(a,y)}{\partial a}(b-a)+\frac{L}{2}|b-a|^{2}$. Applying this $L$-smoothness with $a=\mathbb{E}_{\rho_{\mathbf{x}}}[h(X,z_{j})],b=\mathbb{E}_{\rho_{x\cup\mathbf{x}^{-i}}}[h(X,z_{j})],y=y_{j}$ and taking average over $j\in\{1,2,\ldots,n\}$, we get $B_{F_{0}}(\rho_{x\cup\mathbf{x}^{-i}},\rho_{\mathbf{x}})\leq\frac{L}{2n}\sum_{j=1}^{n}\left|\frac{h(x^{i},z_{j})}{N}\right|^{2}\leq\frac{LR^{2}}{2N^{2}}$. Moreover, we assume $\sup_{|a|\leq R,y\in\mathcal{Y},x\in\mathbb{R}^{d},z\in\mathcal{Z}}\|\partial_{1}\ell(a,y)\partial_{x}h(x,z)\|\leq R^{\prime}$. Then, by the contraction principle (Proposition 5.4.3 in Bakry et al. (2014)) with Theorem 1.4 in Brigati & Pedrotti (2024), Assumption 2 holds with $\alpha=\frac{2\lambda^{\prime}}{\lambda}\exp\left(-\frac{R^{\prime 2}}{2\lambda\lambda^{\prime}}-\frac{4R^{\prime}}{\sqrt{2\lambda\lambda^{\prime}}}\right)$ (see also Lemma 6 in Chewi et al. (2024)).
The defective LSI developed by Chen et al. (2022) is the key condition to study MFLD in finite-particle setting. We here derive an improved variant by exploiting the problem structure and uniform directional LSI. The proof can be found in Appendix A. 1.
@@
Lemma 1 gives an approximation error bound between $\mu^{(N)}$ and $\mu^{\otimes N}_{*}$, which shrinks up to $B/N$ error as $\mu^{(N)}\to\mu^{(N)}_{*}$ and shrinks to zero by additionally taking $N\to\infty$, meaning that each particle of the system $(X^{1},\ldots,X^{N})\sim\mu^{(N)}$ becomes independent to each other. Compared to the original result Chen et al. (2022), the particle approximation term $B/N$ is independent of $\alpha$, similar to Nitanda (2024). Note that whereas Nitanda (2024) only consider the case of $\mu^{(N)}=\mu^{(N)}_{*}$, our result allows for any distribution $\mu^{(N)}$ at the cost of the Fisher information $\mathrm{FI}(\mu^{(N)}\|\mu^{(N)}_{*})$. Lemma 1 can be indeed regarded as an extended LSI on the finite-particle system and nonlinear mean-field objective, where Fisher information is lower bounded by the optimality gap up to $B/N$ error. In particular, when $F_{0}$ is the linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]~{}(\exists f:\mathbb{R}^{d}\to\mathbb{R})$, the lemma reproduces the standard LSI on $\mu^{(N)}_{*}$:
@@
By the standard argument of Langevin dynamics (e.g., Vempala & Wibisono (2019)) and Lemma 1, we get
@@
We note that the particle approximation term $B/N$ in Theorem 1 is independent of LSI-constants, whereas the error $O(\frac{\lambda}{\alpha^{\prime}N})$ obtained in Suzuki et al. (2023a) scales inversely with an LSI constant $\alpha^{\prime}$ on the proximal Gibbs distribution which can be exponentially small as $\lambda\to 0$. Whereas the term $B/N$ is comparable to Nitanda (2024); Chewi et al. (2024), our convergence rate $\exp(-2\alpha\lambda t)$ in optimization is faster since their results rely on the LSI constant $\bar{\alpha}$ on $\mu^{(N)}_{*}$ which is smaller than $\alpha=\frac{\lambda^{\prime}}{\lambda}\exp\left(-\frac{R^{\prime 2}}{\lambda\lambda^{\prime}}-\frac{4R^{\prime}}{\sqrt{\lambda\lambda^{\prime}}}\right)$ (Example 1) in general [^1] However, we note PoC result obtained by Chewi et al. (2024) is applicable to non-bounded activation functions such as ReLU.. For instance, Chewi et al. (2024) estimated the LSI constant $\bar{\alpha}\gtrsim\frac{\lambda^{\prime}}{\lambda}\exp\left(-O\left(\frac{1}{\lambda^{\prime}}+\frac{1}{\lambda\lambda^{\prime}}+\frac{1}{\lambda^{2}\lambda^{\prime 3}}\right)\right)$.
@@ 4.1 Point-wise model approximation error
$\leq\frac{4R^{2}}{N}+8R^{2}\sqrt{\frac{\mathrm{KL}(\mu^{(N)}\|\mu_{*}^{\otimes N})}{N}}.$
@@
For simplicity of comparison, we consider the bound $\frac{4R^{2}}{MN}+\frac{8R^{2}}{M}\sqrt{\frac{B}{\lambda N}}+\frac{2R^{2}B}{\lambda N}$ obtained in the limit $k\to\infty,~{}\eta\to 0$. Then, we see that this bound becomes smaller than the error $\frac{4R^{2}}{N}+8R^{2}\sqrt{\frac{B}{\lambda N}}$ achieved by single network training (Theorem 2) by using sufficiently large $M,N$ becase of the better dependence on $M,N$.
@@
Our result can extend to randomly pruned networks. That is, we consider randomly pruning $(N-s)$-neurons after training the MFNN of $N$-neurons. Then, we get the counterpart of Proposition 1 as follows.
@@
where $\mu^{(N)}_{1:s}$ is the distribution of remaining neurons. Theorem 3 can also extend to the ensemble model of randomly pruned networks in the same way.
@@ 4.2 Uniform model approximation error
For example, $h(x,z)=\frac{R}{3}(\tanh(x^{1\top}z+x^{2})+2\tanh(x^{3})),~{}(x^{1}\in\mathbb{R}^{d},(x^{2},x^{3})\in\mathbb{R}^{2})$, used in Suzuki et al. (2023b) satisfies the above assumption with $\beta=\frac{R}{3}$ due to $1$-Lipschitz continuity of $\tanh$.
Given random variables $\{\mathbf{X}_{j}\}_{j=1}^{M},~{}(\mathbf{X}_{j}=(X^{1}_{j},\ldots,X^{N}_{j}))$, we consider the empirical Rademacher complexity of the function class =$\mathcal{F}=\{x\mapsto h(x,z)\mid z\in\mathcal{Z}\}$:
^$\hat{\mathcal{R}}_{N,M}(\mathcal{F})=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\left|\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\sigma^{i}_{j}f(X^{i}_{j})\right|\right],$
@@
The lemma follows directly by applying the uniform law of large numbers to the function class $\mathcal{F}$ (see, for instance, Mohri et al. (2012). ∎
The complexity $\mathbb{E}_{\{\mathbf{X}_{j}\}_{j=1}^{M}}\left[\hat{\mathcal{R}}_{N,M}(\mathcal{F})\right]$ can be then evaluated by Dudley’s entropy integral (Lemma A) under Assumption 5 and the boundedness $|h(x,z)|\leq R$. By using the variational formulation of KL-divergence (e.g., Corollary 4.15 in Boucheron et al. (2013)), we translate the result of Lemma 3 into the approximation error of an ensemble model obtained by $M$ independent parallel iterates $\mathbf{X}_{j,k}\sim\mu^{(N)}_{k}~{}(j=1,2,\ldots,M)$ of MFLDs. Combining these techniques with Theorem 1, we conclude the following theorem.
@@
Here, the $\tilde{O}$-notation hides logarithmic factors. As for the concrete bound and proofs, see Appendix A. 3. The term $\sqrt{\frac{d}{MN}+\frac{B}{\lambda N}+\frac{d\lambda}{MN(\lambda+MB)}}$ represents the particle approximation error due to finite $N$ particles, and even when $M=1$, they improve upon the bound in Suzuki et al. (2023a, b) by removing the LSI constant $\alpha$ from the corresponding term. And the upper bound shows the improvement as $M$ increases.
@@
Llama2 7B LoRA (best) $79.48$ $82.43$ $81.77$ $80.60$ $67.75$ $80.47$ $70.37$ $86.67$ $78.69$
PoC merge $81.17$ $84.60$ $85.16$ $86.60$ $72.53$ $86.62$ $72.45$ $92.79$ $82.74$
Llama3 8B LoRA (best) $81.22$ $89.50$ $86.74$ $86.00$ $79.86$ $90.53$ $72.91$ $95.34$ $85.26$
PoC merge $82.04$ $89.39$ $89.27$ $89.20$ $83.28$ $92.30$ $76.33$ $96.58$ $87.30$
@@ 5.2 LoRA for finetuning language models
We use commonsense reasoning datasets Hu et al. (2023): SIQA, PIQA, WinoGrande, OBQA, ARC-c, ARC-e, BoolQ, and HellaSwag, and use language models: Llama2-7B Touvron et al. (2023) and Llama3-8B Dubey et al. (2024). We first optimize the multiple LoRA parameters $\{(\mathbf{A}_{j},\mathbf{B}_{j})\}_{j=1}^{M}$ using noisy AdamW where $\sqrt{2\lambda\eta_{k}}\xi_{k}~{}(\textrm{step-size}~{}\eta_{k},~{}\textrm{standard Gaussian noise}~{}\xi_{k})$ is added to each parameter update of AdamW Loshchilov & Hutter (2019). Then, we merge them into $\Delta W=\frac{1}{MN}\sum_{i,j}b^{i}_{j}a^{i\top}_{j}\in\mathbb{R}^{k\times d}$. Hyperparameters are set to $N=32,M=8,\lambda=10^{-5}$ and the number of epochs is $3$. In Table 1, we compare the accuracy of the merged parameter with the base parmeters of LoRA. For LoRA, the table presents the best result achieved among eight different LoRA parameters based on the average accuracy across all datasets. For both models, we observed that PoC-based model merging significantly improves the finetuning performance. We also verify the performance using one-epoch training to examine the effect of Gaussian noise in parameter updates. See Appendix B. 3 for more details.
@@
We established an improved PoC for MFLD that accelerates optimization speed in Nitanda (2024); Chewi et al. (2024) while achieving the same particle complexity $O(1/N)$. We then translated this result into model approximation error bounds, and derived a PoC-based model ensemble method with an empirical verification. Moreover, we substantiated its applicability to fine-tuning language models using LoRA.

### v2

@@ 2.2 Mean-field Langevin dynamics and uniform-in-time propagation of chaos
and $L_{2}$-regularizaton $r(x)=\lambda^{\prime}\|x\|_{2}^{2}$. We assume that $\sup_{x\in\mathbb{R}^{d},z\in\mathcal{Z}}|h(x,z)|\leq R$ and that for any $y\in\mathbb{R}$, $\ell(\cdot,y)$ is convex and $L$-smooth; there exists $L>0$ such that for any $a,b\in\mathbb{R}$, $\ell(b,y)\leq\ell(a,y)+\frac{\partial\ell(a,y)}{\partial a}(b-a)+\frac{L}{2}|b-a|^{2}$. Applying this $L$-smoothness with $a=\mathbb{E}_{\rho_{\mathbf{x}}}[h(X,z_{j})],b=\mathbb{E}_{\rho_{x\cup\mathbf{x}^{-i}}}[h(X,z_{j})],y=y_{j}$ and taking average over $j\in\{1,2,\ldots,n\}$, we get $B_{F_{0}}(\rho_{x\cup\mathbf{x}^{-i}},\rho_{\mathbf{x}})\leq\frac{L}{2n}\sum_{j=1}^{n}\left|\frac{h(x^{i},z_{j})}{N}\right|^{2}\leq\frac{LR^{2}}{2N^{2}}$. Therefore, the Holley-Stroock argument (Holley and Stroock,, 1987) with Eq. (3) implies an LSI constant $\alpha=\alpha_{0}\exp(-\frac{2LR^{2}}{\lambda N})$ converging to $\alpha_{0}$ as $N\to\infty$, where $\alpha_{0}=\frac{2\lambda^{\prime}}{\lambda\exp(O(\lambda^{-1}))}$ is the LSI-constant of the proximal Gibbs distribution $\hat{\rho}_{\mathbf{x}}$.
The following defective entropy sandwich and defective LSI are key results in studying MFLD in the finite-particle setting. The proofs can be found in Appendix A. 1. For $\mathbf{X}\sim\mu^{(N)}$, we denote by $\mu^{(N)}_{i|-i}(\cdot|\mathbf{x}^{-i})$ the conditional distribution of $X^{i}$ conditioned by $\mathbf{X}^{-i}=\mathbf{x}^{-i}$.
Lemma 1 (Defective entropy sandwich).
Suppose Assumption 4 holds. Then, for any $\mu^{(N)}\in\mathcal{P}_{2}(\mathbb{R}^{dN})$,
$\hskip-5.69054pt\frac{\lambda}{N}\mathrm{KL}(\mu^{(N)}\|\mu^{\otimes N}_{*})+\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[B_{F_{0}}(\rho_{\mathbf{X}},\mu_{*})]$
$\hskip-5.69054pt=\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)})-\mathcal{L}(\mu_{*})$
$\hskip-5.69054pt\leq\frac{B}{N}+\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))\right].$
The lemma 1 can be viewed as the finite-particle counterpart of the entropy sandwich established in Nitanda et al., (2022); Chizat, (2022):
[…]
Eq. (3) says that the conditional distribution $\nu_{i|-i}(\cdot|\mathbf{X}^{-i})$ approximates the proximal Gibbs distribution $\hat{\rho}_{\mathbf{X}}$, where $\rho_{\mathbf{X}}$ is an empirical distribution consisting of $\mathbf{X}$. Therefore, we expect $\nu_{i|-i}(\cdot|\mathbf{X}^{-i})$ to play a role analogous to the proximal distribution and lead to an entropy sandwich. In fact, we can confirm that Lemma 1 with $\mu^{(N)}=\mu^{\otimes N}$ reproduces the above entropy sandwich in the infinite-particle system by taking $N\to\infty$ under regular conditions.
The defective LSI was originally established by Chen et al., (2022). We here derive an improved variant built upon the defective entropy sandwich. The proof can be found in Appendix A. 1.
@@
Lemma 2 gives an approximation error bound between $\mu^{(N)}$ and $\mu^{\otimes N}_{*}$, which shrinks up to $B/N$ error as $\mu^{(N)}\to\mu^{(N)}_{*}$ and shrinks to zero by additionally taking $N\to\infty$, meaning that each particle of the system $(X^{1},\ldots,X^{N})\sim\mu^{(N)}$ becomes independent to each other. Compared to the original result Chen et al., (2022), the particle approximation term $B/N$ is independent of $\alpha$, similar to Nitanda, (2024). Note that whereas Nitanda, (2024) only consider the case of $\mu^{(N)}=\mu^{(N)}_{*}$, our result allows for any distribution $\mu^{(N)}$ at the cost of the Fisher information $\mathrm{FI}(\mu^{(N)}\|\mu^{(N)}_{*})$. Lemma 2 can be indeed regarded as an extended LSI on the finite-particle system and nonlinear mean-field objective, where Fisher information is lower bounded by the optimality gap up to $B/N$ error. In particular, when $F_{0}$ is the linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]~(\exists f:\mathbb{R}^{d}\to\mathbb{R})$, the lemma reproduces the standard LSI on $\mu^{(N)}_{*}$:
@@
By the standard argument of Langevin dynamics (e.g., Vempala and Wibisono, (2019)) and Lemma 2, we get
@@
We note that the particle approximation term $B/N$ in Theorem 1 is independent of LSI-constants, whereas the error $O(\frac{\lambda}{\alpha^{\prime}N})$ obtained in Suzuki et al., 2023a scales inversely with an LSI constant $\alpha^{\prime}$ on the proximal Gibbs distribution which can be exponentially small as $\lambda\to 0$. Whereas the term $B/N$ is comparable to Nitanda, (2024); Chewi et al., (2024), our convergence rate $\exp(-2\alpha\lambda t)$ in optimization is faster since their results rely on the LSI constant $\bar{\alpha}$ on $\mu^{(N)}_{*}$ which is smaller than $\alpha_{0}$ in Example 1111However, we note PoC result obtained by Chewi et al., (2024) is applicable to non-bounded activation functions such as ReLU.. For instance, Chewi et al., (2024) estimated the LSI constant $\bar{\alpha}\gtrsim\frac{\lambda^{\prime}}{\lambda}\exp\left(-O\left(\frac{1}{\lambda^{\prime}}+\frac{1}{\lambda\lambda^{\prime}}+\frac{1}{\lambda^{2}\lambda^{\prime 3}}\right)\right)$.
@@ 4.1 Point-wise model approximation error
$\leq\frac{4R^{2}}{N}+8R^{2}\sqrt{\frac{\mathrm{KL}(\mu^{(N)}\|\mu_{*}^{\otimes N})}{N}}.\$
@@
For simplicity, we consider the bound $\frac{4R^{2}}{MN}+\frac{8R^{2}}{M}\sqrt{\frac{B}{\lambda N}}+\frac{2R^{2}B}{\lambda N}$ obtained in the limit $k\to\infty,~\eta\to 0$. This bound indicates that increasing $M$ offers better scalability than increasing $N$, as long as the second term dominates. In fact, under the constraint $MN=\Theta(K)$, where $K$ denotes the total number of neurons, the bound suggests a non-trivial choice for the number of networks $M$. Rewriting the bound using $M$ and $K$, we obtain $O\left(\frac{1}{K}+\frac{1}{\sqrt{\lambda MK}}+\frac{M}{\lambda K}\right)$. This shows that $M$ induces a trade-off, and the bound achieves its minimum value of $O\left(\frac{1}{(\lambda K)^{2/3}}\right)$ when $M=\lambda^{1/3}K^{1/3}$. This phenomenon arises because training multiple networks enhances the independence among particles.
@@
Our result can extend to randomly pruned networks. That is, we consider randomly pruning (Ns)(N-s)-neurons after training the MFNN of $N$-neurons. Then, we get the counterpart of Proposition 1 as follows.
@@
where $\mu^{(N)}_{1:s}$ is the distribution of remaining neurons. Moreover, Theorem 3 can also extend to the ensemble model of randomly pruned networks in the same way.
@@ 4.2 Uniform model approximation error
For example, $h(x,z)=\frac{R}{3}(\tanh(x^{1\top}z+x^{2})+2\tanh(x^{3})),~(x^{1}\in\mathbb{R}^{d},(x^{2},x^{3})\in\mathbb{R}^{2})$, used in Suzuki et al., 2023b satisfies the above assumption with $\beta=\frac{R}{3}$ due to $1$-Lipschitz continuity of $\tanh$.
Given random variables $\{\mathbf{X}_{j}\}_{j=1}^{M},~(\mathbf{X}_{j}=(X^{1}_{j},\ldots,X^{N}_{j}))$, we consider the empirical Rademacher complexity of the function class $\mathcal{F}=\{x\mapsto h(x,z)\mid z\in\mathcal{Z}\}$:
$\hat{\mathcal{R}}_{N,M}(\mathcal{F})=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\left|\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\sigma^{i}_{j}f(X^{i}_{j})\right|\right],$
@@
The lemma follows directly by applying the uniform law of large numbers to the function class $\mathcal{F}$ (see, for instance, Mohri et al., (2012). ∎
The complexity $\mathbb{E}_{\{\mathbf{X}_{j}\}_{j=1}^{M}}\left[\hat{\mathcal{R}}_{N,M}(\mathcal{F})\right]$ can be then evaluated by Dudley’s entropy integral (Lemma A) under Assumption 5 and the boundedness $|h(x,z)|\leq R$. By using the variational formulation of KL-divergence (e.g., Corollary 4.15 in Boucheron et al., (2013)), we translate the result of Lemma 4 into the approximation error of an ensemble model obtained by $M$ independent parallel iterates $\mathbf{X}_{j,k}\sim\mu^{(N)}_{k}~(j=1,2,\ldots,M)$ of MFLDs. Combining these techniques with Theorem 1, we conclude the following theorem.
@@
Here, the $\tilde{O}-$notation hides logarithmic factors. As for the concrete bound and proofs, see Appendix A. 3. The term $\sqrt{\frac{d}{MN}+\frac{B}{\lambda N}+\frac{d\lambda}{MN(\lambda+MB)}}$ represents the particle approximation error due to finite $N$ particles, and even when $M=1$, they improve upon the bound in Suzuki et al., 2023a ; Suzuki et al., 2023b by removing the LSI constant $\alpha$ from the corresponding term. And the upper bound shows the improvement as $M$ increases.
Llama2 7B LoRA (32, best) $79.48$ $82.43$ $81.77$ $80.60$ $67.75$ $80.47$ $70.37$ $86.67$ $78.69$
LoRA (256) $69.95$ $69.69$ $69.61$ $61.40$ $47.44$ $61.15$ $63.73$ $47.27$ $61.28$
PoC merge $81.17$ $84.60$ $85.16$ $86.60$ $72.53$ $86.62$ $72.45$ $92.79$ $82.74$
Llama3 8B LoRA (32, best) $81.22$ $89.50$ $86.74$ $86.00$ $79.86$ $90.53$ $72.91$ $95.34$ $85.26$
LoRA (256) $81.06$ $87.60$ $87.61$ $84.60$ $78.92$ $90.06$ $75.11$ $94.98$ $84.99$
PoC merge $82.04$ $89.39$ $89.27$ $89.20$ $83.28$ $92.30$ $76.33$ $96.58$ $87.30$
@@ 5.2 LoRA for finetuning language models
We use commonsense reasoning datasets Hu et al., (2023): SIQA, PIQA, WinoGrande, OBQA, ARC-c, ARC-e, BoolQ, and HellaSwag, and use language models: Llama2-7B Touvron et al., (2023) and Llama3-8B Dubey et al., (2024). We first optimize the multiple LoRA parameters $\{(\mathbf{A}_{j},\mathbf{B}_{j})\}_{j=1}^{M}$ using noisy AdamW where $\sqrt{2\lambda\eta_{k}}\xi_{k}~(\textrm{step-size}~\eta_{k},~\textrm{standard Gaussian noise}~\xi_{k})$ is added to each parameter update of AdamW Loshchilov and Hutter, (2019). Then, we merge them into $\Delta W=\frac{1}{MN}\sum_{i,j}b^{i}_{j}a^{i\top}_{j}\in\mathbb{R}^{k\times d}$. Hyperparameters are set to $N=32,M=8,\lambda=10^{-5}$ and the number of epochs is $3$. In Table 1, we compare the accuracy of the merged parameter with the base parmeters of LoRA. For LoRA, the row “LoRA (32, best)” presents the best result achieved among eight different LoRA parameters based on the average accuracy across all datasets. For both models, we observed that PoC-based model merging significantly improves the finetuning performance. For comparison under the same computational budget $MN=256$, we report the results obtained by LoRA with $M=1$ and $N=256$ in the row “LoRA (256)”. We also verify the performance using one-epoch training to examine the effect of Gaussian noise in parameter updates. See Appendix B. 3 for more details.
@@
We established an improved PoC for MFLD that accelerates optimization speed in Nitanda, (2024); Chewi et al., (2024) while achieving the same particle complexity $O(1/N)$. We then translated this result into model approximation error bounds, and derived a PoC-based model ensemble method with an empirical verification. Moreover, we substantiated its applicability to fine-tuning language models using LoRA.
@@
Acknowledgements
This research is supported by the National Research Foundation, Singapore, Infocomm Media Development Authority under its Trust Tech Funding Initiative, and the Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-004). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, Infocomm Media Development Authority, and the Ministry of Digital Development and Information. TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2115).